Suppose that permanent income is calculated as the average of income over the past 5 years; that is,
YP=1/5(Y+Y_1+Y_2+Y_3+Y_4)
Suppose further that consumption is given as C=0.9YP
I).If you have earn 20,000 per year for the past 10 years, what is your permanent income?
1
Expert's answer
2021-09-07T19:14:33-0400

"YP \ = \frac{1}{5}(Y + Y\_1 + Y\_2 + Y\_3+ Y\_4)"


I)

"YP = \frac{1}{5} \times (20000 \times 5)"

"YP = 20000"